nolegirl24:

Which exponential function goes through the points (1, 32) and (4, 2048)? f(x) = 4(4)x f(x) = 8(2)−x f(x) = 8(4)x f(x) = 4(4)−x

Vocaloid:

well using process of elimination the exponent is probably not negative considering that the function increases so much

nolegirl24:

So it is A or B bc the others are -

Vocaloid:

let x = 1 and see which function equals 32

Vocaloid:

8(4)^1 = ?

nolegirl24:

Sorry im back

nolegirl24:

I got 4096

Vocaloid:

as a hint, anything raised to an exponent of 1 is just itself so 4^1 is just 4

Vocaloid:

8(4^1) = 8*4 = 32 making it our answer (C)

nolegirl24:

Ohhh dang it!!!!